A random sample of 16 values from a normal population showed a mean 41.5 inches and sum of squares of deviations from this mean equal to 135 square inches. Show that assumption of a mean of 43.5 inches for the population is not reasonable. Obtain 95% and 99% confidence limits for same.


1
Expert's answer
2021-11-09T16:12:03-0500

"\bar{x} =41.5 \\

s = \sqrt{\frac{135}{16-1}}=3 \\

n=16"

Confidence interval:

"CI = \bar{x} ±" Margin of error (E)

Standard error "sx = \frac{s}{\sqrt{n}}"

df= n-1=15

Critical value for 95% confidence interval "t_c=2.49"

"CI = 41.5 ± 2.490 \times 0.75 \\

CI = 41.5 ± 1.8674 \\

CI =(39.632, 43.367)"

Critical value for 99% confidence interval "t_c=3.286"

"CI = 41.5 ± 3.286 \times 0.75 \\

CI = 41.5 ± 2.4645 \\

CI =(39.035, 43.964)"

The assumption of a mean of 43.5 is NOT reasonable at 95% confidence interval as it does not fall within the interval range.

The assumption of a mean of 43.5 is reasonable at 99% confidence interval as it falls within the interval range.